OpenStudy (anonymous):

find the point of intersection of the lines 2x-3y=-2 and 9x+12y=25

OpenStudy (anonymous):

I know you use matrices but what do you do after you plug it in the calculator

OpenStudy (anonymous):

you don't have to explicitly use matrices you can solve by substitution or elimination...

OpenStudy (anonymous):

think about multiplying the first equation by 4 and then adding it to the second equation... see what you get

OpenStudy (anonymous):

yeah i know those but my teacher wants matrices haha, the answer would x= -59 and y = 47

OpenStudy (anonymous):

and i dont know how to work it with matrices

OpenStudy (anonymous):

on the calculator

OpenStudy (anonymous):

\[\left[\begin{matrix}2 & -3 \\ 9 & 12\end{matrix}\right] \]

OpenStudy (anonymous):

use the calculator to find the inverse, multiply it by \[\left(\begin{matrix}-2 \\ 25\end{matrix}\right)\]

OpenStudy (anonymous):

ok i did and it comes out as 1 and 1.333333333333

OpenStudy (anonymous):

yep